Problem 1 Solve y + 4y 1, if 0<t<T, y(0) = 0, y() = 0. if <t<oo.
Answer #1

Case for T s t < 0 y+ 4y zo S.E is (0²+4 ) yao Ao is D²+ 4=0 D= +2 solution is y = Ca Cosatt cy Sinat Since Now so y (t) is4 > - I Cos 2T + 1 = (3 con 25+ Cu Sin 27 - 1 / cat o > | C3=0 th 1/ sto - Now - T Son y (t) = +2 Sinot 4 for Tetc y (t) =